algebraic fractions 5/8 - 3/10 =
Answer by mykailao(6)

5/8 - 3/10
Make their denominators equal first. Their LCM (least common multiple) is 40, so 40 is going to be the denominator.
5/8 = 25/40 (40/8 is 5. 5x5=25)
3/10 = 12/40 (40/10 is 4. 4x3=12)
You can now subtract because they have the same denominator. 25-12 = 13 is the numerator.
Therefore the answer is 13/40.